OpenStudy (anonymous):

What is the intercept of the line passing through (5, -6) with a slope of -1/7 ?

OpenStudy (anonymous):

First we need to generate the equation of the line Y+6=-1/7(x-5) Simplify: X+7y+37=0

OpenStudy (anonymous):

X intercept is the value pf x when y is 0 X=-37 Y intercept is the value of y when x is 0 Y=-37/7